OpenStudy (anonymous):

The polynomial x 3 + 5x 2 -57x -189 expresses the volume, in cubic inches, of a shipping box, and the width is (x+3) in. If the width of the box is 15 in., what are the other two dimensions? ( Hint: The height is greater than the depth.)

OpenStudy (anonymous):

A. height: 19 in. depth: 5 in. B. height: 21 in. depth: 5 in. C. height: 19 in. depth: 7 in.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@DecentNabeel

OpenStudy (anonymous):

easy

OpenStudy (anonymous):

Can you please help me? I've already answered it wrong once lol @timo86m

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

maybe polynomial long division will work here. Divide your polynomial by x+3 maybe.

OpenStudy (anonymous):

But how do i even get to that part? @timo86m

OpenStudy (anonymous):

because usually u have to factor. And you can take x+3 as a hint that it is one of the factors.

OpenStudy (anonymous):

This is the factor that i figured out (x+3)(x−7)(x+9)

OpenStudy (anonymous):

@timo86m

OpenStudy (anonymous):

here is how i picture it in my head You are given x+3. U will need 2 other similar expressions to get a volume call them x+a x+b x+3 <- given And together they form (x+a) * (x+b) * (x+3) = Volume height*depth* width = volume

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

Where do i go from there? @timo86m

OpenStudy (anonymous):

width = 15 so you are told x+3=15 x=15-3 x=12 therefor x = 12 Plug it in :)

OpenStudy (anonymous):

(x+3)(x−7)(x+9) X=12

OpenStudy (anonymous):

( Hint: The height is greater than the depth.)

OpenStudy (anonymous):

 find two other sides

OpenStudy (anonymous):

Is it A?

OpenStudy (anonymous):

@timo86m

OpenStudy (anonymous):

not sure yet

OpenStudy (anonymous):

Thank you @timo86m

OpenStudy (anonymous):

x+9=12+9=21 x-7=12-7=5